OpenStudy (anonymous):

use implicit differnetiation to solve 72(sqrtx) + 12(sqrty)= 36

OpenStudy (anonymous):

\[72\sqrt{x}+12\sqrt{y}=36\Rightarrow 72x^{\frac{ 1 }{ 2 }}+12y^{\frac{ 1 }{ 2 }}=36\]

OpenStudy (anonymous):

i got -3rooty/rootx ?? does that sound right?

OpenStudy (anonymous):

\[\Rightarrow 36x^{-\frac{ 1 }{ 2 }}\cdot \frac{ dx }{ dx } + 6y^{-\frac{ 1 }{ 2 }}\cdot \frac{ dy }{ dx } = 0\]

OpenStudy (anonymous):

nope,\[\frac{ dy }{ dx } = -\frac{ 36\sqrt{y} }{ 6\sqrt{x} }=-\frac{ 6\sqrt{y} }{ \sqrt{x} }\]